# AL-Bench: A Benchmark for Automatic Logging

### v2

@@ 4.4. RQ3: How well do the generated log statements print logs in dynamic evaluation?
Method Cosine Similarity BLEU BLEU-1 BLEU-4 ROUGE-1 ROUGE-L
$FastLog$ 0.213 16.172 24.829 15.963 24.427 23.841
$UniLog_{{cl}}$ 0.174 13.625 19.015 13.423 19.167 18.749
$UniLog_{{ds}}$ 0.130 11.607 14.674 11.556 14.002 13.869
$LANCE$ 0.099 8.201 11.342 7.971 11.235 11.033
$LEONID_{M}$ 0.072 5.466 7.946 5.303 8.095 7.944
$LEONID_{S}$ 0.044 3.220 4.908 2.955 5.099 5.012
Table 6 compares the semantic similarity between logs from source and predicted log statements. The results indicate that logs from the predicted statements significantly deviate from the ground truth, with consistently low scores across metrics like Cosine Similarity, BLEU, and ROUGE. Specifically, low similarity scores indicate that generated logs frequently fall short of matching expected outputs. Limiting each code snippet to a single log statement often sharply compromises log similarity, especially when multiple statements are needed. This limitation is particularly clear when comparing $LEONID_{S}$ and $LEONID_{M}$. Although $LEONID$ overall performs poorly, $LEONID_{M}$ stands out as the only tool capable of generating multiple log statements, which enables it to outperform $LEONID_{S}$ in similarity scores. This difference underscores the importance of tools being able to determine the appropriate number of log statements for accurate and effective logging, rather than assuming a single statement suffices.
Table 7. False Positive and False Negative Logging Rates across Methods. FPLG means the predicted log statement record logs when it should not, and FNLG means the predicted log statement does not record logs when it should.
Method FPLG Rate FNLG Rate
@@
$LEONID_{M}$ 7.32% 11.6%
Table 8. FPLG and FNLG Reason Analysis. For FNLG, the major reason is beyond the execution path, and for FPLG, the major reason is lower verbosity level.
Situation Reason Number
FNLG Beyond Execution Path 35
@@
FPLG Higher Verbosity Level 30
@@
To understand the low semantic similarity scores, we examined the log generation process and found many instances where predicted logs were either redundant or missed key information present in the original logs. We quantify this issue using two metrics: False Positive Log Generation (FPLG) and False Negative Log Generation (FNLG), as reported in Table 8. For example, FastLog reports a 9.2% False Positive Log Rate (FPLR), meaning logs record redundant information in 9.2% of cases, and an 18.28% False Negative Log Rate (FNLR), indicating expected information in logs were missing in 18.28% of cases. We sampled 100 examples from FastLog, the state-of-the-art (SOTA) model, and manually analyzed the reasons for mismatches with the original logs. The results are presented in Table 8. We found that the primary reasons for failures in FNLG and FPLG differ significantly. For FNLG, the most common issue was the predicted log statements being beyond the execution path (35 cases), followed by lower verbosity levels (24 cases), and a smaller number caused by wrong code format (4 cases). In contrast, for FPLG, the main problem was higher verbosity levels (30 cases), with a few instances of log statements being beyond the execution path (3 cases). Overall, verbosity mismatches and execution path discrepancies were the dominant contributors, highlighting challenges in aligning predicted logs with actual logging requirements. The factors leading to FPLG and FNLG underscore a critical issue: while static metrics offer valuable insights into the quality of generated log statements, the actual logs are shaped by numerous contextual factors. Without a thorough understanding of the execution context, it is not possible to comprehensively evaluate the quality of log statements. Even minor discrepancies can cause significant deviations between generated logs and source logs. For instance, while the predicted log statement may capture the key information required to reflect system behavior, its effectiveness can be compromised if it is not positioned along a critical execution path or if its verbosity level is mismatched. In such cases, the log statement may fail to record essential information when key events occur. This issue majorly arises from the tools’ lack of awareness of verbosity thresholds and the control graph of the code, which limits their ability to adjust verbosity and determine appropriate log positions based on the context or execution requirements. These two limitations highlight that current logging tools lack the adaptability and context-awareness needed for effective real-world application.
@@
Answer to RQ3. The best predicted log statements by FastLog achieve only 0.231 cosine similarity with the original logs. Many predictions record redundant information, while others miss key details. The missing key information is primarily due to the prediction being placed beyond the execution path during important events, while setting a higher verbosity level in the log statements leads to redundancy. This result highlights that automatic logging tools still have significant room for improvement.
@@ 5. Threats to Vadility
  • Construct Validity: We use BLEU to assess the quality of the generated log messages. Although text similarity metrics may not fully capture the quality of the generated text (Gros et al., 2020; Roy et al., 2021), we follow previous works (Xie et al., 2023; Xu et al., 2024; He et al., 2018a; Ding et al., 2023) to use BLEU as widely-accepted quality measures for the generated log messages. We also adapt Cosine Similarity based on TF-IDF to evaluate the quality of log content. Since log content, excluding the log header, is primarily composed of natural language, cosine similarity effectively captures the semantic meaning (Salton et al., 1975).
  • Internal Validity: The second threat to validity concerns reproducing the baseline. To minimize inconsistencies, we adapted the released models from previous work (Xie et al., 2023; Mastropaolo et al., 2022, 2023) and sought guidance from the authors of the closed-source model UniLog (Xu et al., 2024) to reproduce the tool under supervision.
  • External Validity: In this work, we evaluate the tools with 10 widely recognized Java projects in different domains. The results of each tool are examined in different requirements of logging practices. The novel evaluation method we propose is not limited to Java projects, we will explore its effects in other programming languages in the future.
@@ 6. Conclusion
This paper introduces AL-Bench, designed specifically for automatic logging tools. It includes a high-quality dataset and a novel dynamic evaluation method focused on runtime logs, addressing key limitations of prior studies and bridging the gap between real-world requirements and existing evaluation frameworks. This dynamic evaluation method assesses both the compilability of predicted log statements and their effectiveness in generating runtime logs. Using AL-Bench, we evaluate popular end-to-end automatic logging tools and find that generated log statements fail to compile in 20.1% to 83.6% of cases. Even the best predictions achieve only 0.213 cosine similarity between generated and ground-truth runtime logs. These results show that automatic logging still has a long way to go.

### v3

@@ 4.4. RQ3: How consistent are runtime logs from auto-logged programs with oracle-generated logs?
Method COS(%) BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
$FastLog$ 21.32 24.83  19.62  17.49 15.96 24.43  18.74 23.84
$UniLog_{{cl}}$ 17.38 19.02  15.62  14.34 13.42 19.17  15.37 18.75
$UniLog_{{ds}}$ 13.04 14.67  12.72  12.02 11.56 14.00  12.23 13.87
$LANCE$ 9.93 11.34  9.21  8.50 7.97 11.23  8.94 11.03
$LEONID_{M}$ 7.19 7.95  6.51  5.87 5.30 8.10  6.43 7.94
$LEONID_{S}$ 4.45 4.91  3.82  2.95 2.96 5.10  3.86 5.01
Table 6 compares the semantic similarity between logs from source and predicted log statements. The results indicate that logs from the predicted statements significantly deviate from the ground truth, with consistently low scores across metrics like Cosine Similarity, BLEU, and ROUGE. Limiting each code snippet to a single log statement often sharply compromises log similarity, especially when multiple statements are needed. This limitation is particularly clear when comparing $LEONID_{S}$ and $LEONID_{M}$. Although $LEONID$ overall performs poorly, $LEONID_{M}$ stands out as the only tool capable of generating multiple log statements, which enables it to outperform $LEONID_{S}$ in similarity scores. This difference underscores the importance of tools being able to determine the appropriate number of log statements for accurate and effective logging, rather than assuming a single statement suffices.
Table 7. False Positive and False Negative Logging Rates across Methods. FPLR means the predicted log statement record logs when it should not, and FNLR means the predicted log statement does not record logs when it should.
Method FPLR FNLR
@@
$LEONID_{M}$ 7.32% 11.60%
Table 8. FPLR and FNLR Reason Analysis. For FNLR, the major reason is beyond the execution path, and for FPLR, the major reason is lower verbosity level.
Situation Reason # False
FNLR Beyond Execution Path 35
@@
FPLR Higher Verbosity Level 30
@@
To understand the low semantic similarity scores, we examined the log generation process and found many instances where predicted logs were either redundant or missed key information present in the original logs. We quantify this issue using two metrics: FPLR and FNLR, as reported in Table 8. For example, FastLog reports a 9.28% FPLR, meaning logs record redundant information in 9.28% of cases, and an 18.28% FNLR, indicating expected information in logs was missing in 18.28% of cases. We sampled 100 examples from FastLog, the state-of-the-art (SOTA) model, and manually analyzed the reasons for mismatches with the original logs. The results are presented in Table 8. We found that the primary reasons for failures in FNLR and FPLR differ significantly. For FNLR, the most common issue was the predicted log statements being beyond the execution path (35 cases), followed by lower verbosity levels (24 cases), and a smaller number caused by wrong code format (4 cases). In contrast, for FPLR, the main problem was higher verbosity levels (30 cases), with a few instances of log statements being beyond the execution path (3 cases). Overall, verbosity mismatches and execution path discrepancies were the dominant contributors, highlighting challenges in aligning predicted logs with actual logging requirements. The factors leading to FPLR and FNLR underscore a critical issue: while static metrics offer valuable insights into the quality of generated log statements, the actual logs are shaped by numerous contextual factors. Without a thorough understanding of the execution context, it is not possible to comprehensively evaluate the quality of log statements. Even minor discrepancies can cause significant deviations between generated logs and source logs. For instance, while the predicted log statement may capture the key information required to reflect system behavior, its effectiveness can be compromised if it is not positioned along a critical execution path or if its verbosity level is mismatched. In such cases, the log statement may fail to record essential information when key events occur. This issue majorly arises from the tools’ lack of awareness of verbosity thresholds and the control graph of the code, which limits their ability to adjust verbosity and determine appropriate log positions based on the context or execution requirements. These two limitations highlight that current logging tools lack the adaptability and context-awareness needed for effective real-world applications.
@@
Answer to RQ3. The best-predicted log statements by FastLog achieve only 21.32% cosine similarity with the original logs. Many predictions record redundant information, while others miss key details. The missing key information is primarily due to the prediction being placed beyond the execution path during important events, while setting a higher verbosity level in the log statements leads to redundancy. This result highlights that automatic logging tools still have significant room for improvement.
@@ 5. Threats to Vadility
  • Construct Validity: The first threat comes from the use of two widely-accepted syntactic metrics in previous logging research (BLEU and ROUGE) (Xie et al., 2024; Xu et al., 2024a; He et al., 2018a; Ding et al., 2023), as they may not fully capture the semantics of the log messages (Gros et al., 2020; Roy et al., 2021). To mitigate this threat, we also adapt Cosine Similarity to evaluate the semantic quality of log messages. Since all log headers are excluded, leaving log messages primarily in natural language, Cosine Similarity can effectively capture the log’s semantic meaning (Salton et al., 1975).
  • Internal Validity: The second threat to validity concerns reproducing the baseline. To minimize inconsistencies, we adapted the released models from previous work (Xie et al., 2024; Mastropaolo et al., 2022, 2023) and sought guidance from the authors of the closed-source model UniLog (Xu et al., 2024a) to reproduce the tool under supervision.
  • External Validity: The third threat comes from the potential data contamination risks in our dataset. To mitigate it, we follow the methodology used by GPT-3 (Brown et al., 2020) to assess data contamination severity, i.e., considering test samples with more than a 13-gram overlap with the original code samples as contaminated. Specifically, we perform a 13-gram overlap analysis between test and training samples of all evaluation methods. Results indicate a peak contamination rate of 1.24% across all training datasets, empirically demonstrating no evidence of data leakage in our framework.
@@ 6. Conclusion
In this paper, we introduced AL-Bench, a comprehensive benchmark for automatic logging tools evaluation, featuring a diverse dataset from 10 established projects and a novel dynamic evaluation methodology. Our benchmark revealed significant limitations in existing tools through both static and dynamic perspectives: substantial accuracy drops in log position, level, and message prediction compared to reported results, high compilation failure rates (20.1%-83.6%), and low runtime log similarity (maximum 21.32%). These findings highlight critical gaps between current tool capabilities and practical requirements. We believe AL-Bench establishes a solid foundation for advancing automatic logging research by providing standardized evaluation metrics and exposing areas requiring improvement.